OpenStudy (anonymous):

Solve the compound inequality 30 < 10u < 120. 3 < u < 4 3 > u > 12 3 > u > 4 3 < u < 12

OpenStudy (anonymous):

Remember we want to get "u" all by itself. Since we have 10u in the middle, dividing the entire inequality by 10 will give us u all by itself. Since 10 is positive we don't need to switch any inequality signs. So, after you divide everything by 10 you'll have your answer

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

no, it wouldn't be A, because 120 divided by 10 isn't 4. When we divide the entire thing by 10 it'll look like this: \[\frac{ 30 }{ 10 } < \frac{ 10u }{ 10 } < \frac{ 120 }{ 10 }\]

OpenStudy (anonymous):

oh so D

OpenStudy (anonymous):

Yes, exactly.

OpenStudy (anonymous):

thxs!

OpenStudy (anonymous):

You're welcome, glad I could help